OpenStudy (anonymous):

Find the sum 1/7 + 3/7 + 3^2 / 7 + 3^3/7 + ..... 3^(n-1) 7

OpenStudy (anonymous):

Hey take 1/7 common out of all terms... you will notice a geometric progression formed...

OpenStudy (anonymous):

\[\frac{1}{7} (3^0 + 3^1 + 3^2 ..... 3^{(n-1)}) \] I hope you are familiar with the sum of geometric progression. \[\frac{a(1-r^n)}{1-r} \] Take 1 as a ; 3 as r

OpenStudy (anonymous):

It says that the answer its \[-\frac{ 1 }{ 14 } (1-3^{n})\]

OpenStudy (anonymous):

Did you try solving it?

OpenStudy (anonymous):

Yes the answer is 4

OpenStudy (anonymous):

Wait! Answer is 4 ? , How did you get it ?

OpenStudy (anonymous):

\[\frac{ 1(1-3^{2}) }{ 1-3 } = \frac{ -8 }{ -2 } = 4\]

OpenStudy (anonymous):

According to the question, you don't know the value of n , so, you can't just put n as 2 ! :P :) Moreover , you missed out that 1/7 :P

OpenStudy (anonymous):

Im lost :( sorry. So it will be \[\frac{ 1(1-3^{n}) }{ 1-3 }\] What I do after

OpenStudy (anonymous):

It's okay to be lost! :D And, what you did, isn't wrong, it's just incomplete. it would be : \[\frac{1}{7}\frac{1(1−3^n)}{1−r}\]

OpenStudy (anonymous):

Now I get it! Your awersome :)

OpenStudy (anonymous):

Haha. :D Thank You! :D